OpenStudy (anonymous):

How Do I Simplify The (-5SQRT8)(-3SQRT5) ?

OpenStudy (saifoo.khan):

\[30\sqrt{10}\]

OpenStudy (anonymous):

How'd You Get That ?

OpenStudy (saifoo.khan):

(-5SQRT8)(-3SQRT5)\[-10\sqrt{2}*-3\sqrt{5}\]\[(-10*-3)(\sqrt{2}*\sqrt{5})\]\[30\sqrt{10}\]

OpenStudy (anonymous):

Where Did The 10SQRT2 Come From ?

OpenStudy (saifoo.khan):

\[-5\sqrt{4*2}\]you know what sqrt of 4 is 2. so\[-5\sqrt{4}\sqrt{2}\]\[-5*2\sqrt{2}\]\[-10\sqrt{2}\]